OpenStudy (anonymous):

For each direct variation, find the constant of variation. then find the value of y when x = 4. 1) y=3 when x = -2

OpenStudy (anonymous):

If y varies directly with x, this means, \[y=kx\]Since they provide x=-2, y=3, we solve for k:\[3=-2k\] \[k=\frac{-3}{2}\]

OpenStudy (anonymous):

To find y when x=4, solve:\[y=\frac{-3}{2}(4)\]This yields,\[y=-6\]

OpenStudy (anonymous):

Really dumb question.......how did you get \[k = -2/3\] from 3=−2k?

OpenStudy (anonymous):

He didn't. He got k=-3/2. From 3=-2k divide both sides by -2 -3/2=k

OpenStudy (anonymous):

That would explain it, thanks.